# countnullcountdimennullnullnulldimennullnullnulldimennullnullnulldimennullnullnulldimen

### v1

@@ Appendix E Minimax analysis for $\mathcal{P}_{2}(\mu,d)$ when $\mu\leq d<2\mu$
$\lim_{m\rightarrow\infty}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\sup_{D\in\mathcal{D}_{m}}\ \,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{m}$ $\geq\lim_{m\rightarrow\infty}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\sup_{p_{m}}\ \,\frac{p_{m}\cdot\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p_{m})}{m}\geq\mu-\frac{d}{2}+\xi$ (E.1)
@@
As we are considering two-point distributions in an i.i.d. setting, the choice of $\mathbb{P}^{m}$ can be characterized by the choice of $\alpha=\alpha(m)$ in (2.10). [^7] If the infimum is not attained we take a sequence $(\alpha(m))_{m\in\mathbb{N}}$ close enough to attaining the infimum, i.e., for which $\sup_{p}\ \,p\cdot\mathbb{P}^{m}_{\alpha(m)}(\sum_{i=1}^{m}X_{i}\geq p)\leq\delta+\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)}\sup_{p}\ \,p\cdot\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p)$ for an arbitrary $\delta>0$. We write $\mathbb{P}^{m}=\mathbb{P}^{m}_{\alpha(m)}$ to emphasize this, and focus on bounding
$\lim_{m\rightarrow\infty}\sup_{p_{m}}\ \,\frac{p_{m}\cdot\mathbb{P}^{m}_{\alpha(m)}(\sum_{i=1}^{m}X_{i}\geq p_{m})}{m}.$ (E.2)
@@
The Chebyshev bound in Lemma A.1 then tells us that
@@
$1-\left(\frac{d}{2\gamma\mu}\right)^{2}\tau_{0}\geq 0.99.$ (E.4)
@@
where the final two inequalities use the choice of $\gamma$ in (E.3) and $\tau_{0}$ in (E.4), respectively. This means that for $\lambda\in(0,\tau_{0}]$, the seller can guarantee a revenue of at least $d/2\geq\mu-\frac{d}{2}+\xi_{0}$.
@@
$\frac{[(m-1)x(\alpha)+y(\alpha)]\cdot\mathbb{P}^{m}_{\alpha(m)}(\sum_{i=1}^{m}X_{i}\geq(m-1)x(\alpha)+y(\alpha))}{m}$ $\approx(1-\alpha^{m})\left(\mu-\frac{d}{2}\right)+\frac{(1-\alpha^{m})}{m(1-\alpha)}\frac{d}{2}.$ (E.5)

### v2

@@ Appendix A Minimax analysis for $\mathcal{P}_{2}(\mu,d)$ when $\mu\leq d<2\mu$
$\lim_{m\rightarrow\infty}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\sup_{D\in\mathcal{D}_{m}}\ \,\frac{\mathrm{REV}(D,\mathbb{P}^{m})}{m}$ $\geq\lim_{m\rightarrow\infty}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\sup_{p_{m}}\ \,\frac{p_{m}\cdot\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p_{m})}{m}\geq\mu-\frac{d}{2}+\xi$ (A.1)
@@
As we are considering two-point distributions in an i.i.d. setting, the choice of $\mathbb{P}^{m}$ can be characterized by the choice of $\alpha=\alpha(m)$ in (2.6). Note that if the infimum is not attained we take a sequence $(\alpha(m))_{m\in\mathbb{N}}$ close enough to attaining the infimum, i.e., for which $\sup_{p}\ \,p\cdot\mathbb{P}^{m}_{\alpha(m)}(\sum_{i=1}^{m}X_{i}\geq p)\leq\delta+\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)}\sup_{p}\ \,p\cdot\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p)$ for an arbitrary $\delta>0$. We write $\mathbb{P}^{m}=\mathbb{P}^{m}_{\alpha(m)}$ to emphasize this, and focus on bounding
$\lim_{m\rightarrow\infty}\sup_{p_{m}}\ \,\frac{p_{m}\cdot\mathbb{P}^{m}_{\alpha(m)}(\sum_{i=1}^{m}X_{i}\geq p_{m})}{m}.$ (A.2)
@@
The Chebyshev bound in Lemma 5.1 then tells us that
@@
$1-\left(\frac{d}{2\gamma\mu}\right)^{2}\tau_{0}\geq 0.99.$ (A.4)
@@
where the final two inequalities use the choice of $\gamma$ in (A.3) and $\tau_{0}$ in (A.4), respectively. This means that for $\lambda\in(0,\tau_{0}]$, the seller can guarantee a revenue of at least $d/2\geq\mu-\frac{d}{2}+\xi_{0}$.
@@
$\frac{[(m-1)x(\alpha)+y(\alpha)]\cdot\mathbb{P}^{m}_{\alpha(m)}(\sum_{i=1}^{m}X_{i}\geq(m-1)x(\alpha)+y(\alpha))}{m}$ $\approx(1-\alpha^{m})\left(\mu-\frac{d}{2}\right)+\frac{(1-\alpha^{m})}{m(1-\alpha)}\frac{d}{2}.$ (A.5)